# Alleviating the Hubble tension with Torsion Condensation (TorC)

### v1

@@ I   Introduction
[…]
@@
$+12\tensor{\mathcal{R}}{{}_{\mu\nu}}\big{(}\tensor{\mathcal{R}}{{}^{[\mu\nu]}}-\tensor{\mathcal{R}}{{}^{\mu\nu}}\big{)}$
@@
[…]
@@
To explore the cosmology of TorC, an equivalent scalar-tensor formulation is introduced. Among the various gravity theories, those based on scalar fields within the framework of modified gravity are one of the most extensively explored models. Scalar-tensor theories involve the coupling of different scalar fields to the metric in a curved but torsion-free spacetime [25]. Note that, in the absence of torsion, the non-Riemannian curvature $\tensor{\mathcal{R}}{}$ becomes the usual Riemannian curvature $\tensor{R}{}$ familiar from GR — this matter is explained in detail in Appendix A. In the cosmological context, scalar fields offer the advantage of inducing accelerated expansion without necessarily violating the isotropy of the Universe, [^3] This is in contrast to vector models, which always ‘point somewhere’; the condensation of a vector field to finite vacuum expectation value is generally understood to constitute an æther [26, 27]. making them compelling candidates for dark energy and early-Universe inflationary mechanisms. It has been shown that in a flat, homogeneous and isotropic Universe, the Lagrangian of PGT can be mapped to a bi-scalar-tensor theory [21]. For TorC, the scalar-tensor equivalent Lagrangian density is
@@
In this work, the cosmological implications of the TorC model are investigated, with particular attention given to its potential to alleviate the Hubble tension. TorC introduces two parameters beyond those in $\Lambda\mathrm{CDM}$: the early-Universe value of the torsion scalar field $\varpi_{r}$ and its velocity, which may be ‘packaged’ into the bare dark energy density parameter $\Omega_{L}$. To constrain the TorC cosmological parameters, the nested sampling algorithm PolyChord [31] is employed, interfaced via Cobaya [32] with cosmological likelihoods and a modified version of CAMB [33]. A combined analysis is carried out between the Planck 2018 CMB data [30] and the 2020 SH0ES supernova measurement [29], assessing the degree of tension between the two datasets using the $R$-statistic [34, 35, 36].
@@ II   Cosmology
 $\Omega_{L}$ Bare dark energy density parameter
 $\varpi_{r}$ Torsion scalar field value in the early Universe
@@
$\Omega_{b}$ Baryon density parameter
$\Omega_{c}$ Cold dark matter density parameter
@@
$A_{s}$ Amplitude of primordial scalar power spectrum
$n_{s}$ Primordial scalar spectral index
@@
This section focuses on the two additional parameters introduced by the model: $\Omega_{L}$, the bare dark energy density parameter, and $\varpi_{r}$, specifying the value of the torsion scalar field in the early Universe, the list of cosmological parameters for TorC is presented in Table 2. In this section, their origin will be outlined: $\Omega_{L}$ arises from the field equations of TorC, while $\varpi_{r}$ is set as an initial condition required for solving them. TorC can be treated as an extension of $\Lambda\mathrm{CDM}$ by absorbing the effects of $\Omega_{L}$ and $\varpi_{r}$ into an effective dark energy density and pressure. This reformulated dark energy is incorporated into a modified version of CAMB to investigate their impact on the CMB power spectrum.
II.1 TorC cosmological parameters: $\Omega_{L}$ and $\varpi_{r}$
@@
$\rho$ $=3M_{\mathrm{P}}^{2}H_{0}^{2}(\Omega_{r}+\Omega_{m}+\Omega_{L}),$ (8a)
$P$ $=3M_{\mathrm{P}}^{2}H_{0}^{2}\left(\frac{\Omega_{r}}{3}-\Omega_{L}\right),$ (8b)
where $H_{0}$ is the Hubble parameter today, $\Omega_{r}$ and $\Omega_{m}$ are the density parameters of radiation and matter, and $\Omega_{L}$ represents the ‘intrinsic’ dark energy density parameter, or the ‘bare’ dark energy density parameter.
@@
$H^{2}=$ $H_{0}^{2}\frac{\Omega_{r}a^{-4}+\Omega_{m}a^{-3}+\Omega_{L}}{\varpi^{2}}$
@@
$H^{2}+\dot{H}^{2}=$ $-H_{0}^{2}\frac{\Omega_{r}a^{-4}+\frac{1}{2}\Omega_{m}a^{-3}-\Omega_{L}}{\varpi^{2}}$
@@
ϖ¨=¨italic-$\ddot{\varpi}=$ $\frac{1}{\left(\varpi^{2}-1\right)\left(3\varpi^{2}+1\right)}$
×$\times\Bigl{(}-3\varpi\left(\varpi^{2}-1\right)^{2}\left(2H^{2}+\dot{H}\right)$
@@
$H^{2}=H_{0}^{2}\left(\Omega_{r}a^{-4}+\Omega_{m}a^{-3}+\Omega_{\Lambda}a^{-4}\right),$ (11)
where $\Omega_{\Lambda}$ is the dark energy density parameter in $\Lambda\mathrm{CDM}$. It is evident that the equations Eq. 9a and Eq. 11 differ in two key ways: first, the density parameters are effectively divided by $\varpi^{2}$ in Eq. 9a; second, an additional term appears in Eq. 9a, involving the scale factor, torsion scalar field, and their derivatives. In standard $\Lambda\mathrm{CDM}$, the dark energy parameter $\Omega_{\Lambda}$ is a constant that can be determined from the density parameters of the other components at $a=1$ and $H=H_{0}$, giving rise to the intuition that the non-gravitational energy budget of the Universe sums to unity when normalized. However, this relationship does not hold for TorC due to the presence of the extra term, similar to how a substantial curvature density would modify the Friedmann equations in GR. As a result, $\Omega_{L}$ cannot be inferred from the other density parameters: it becomes an independent parameter of the model.
@@
using power series initial conditions for $\varpi$ and $a$ in the early Universe. As the bare dark energy density parameter $\Omega_{L}$ plays a crucial role in shaping the late-time expansion of the Universe, it leads to a modified value of the current Hubble parameter $H_{0}$. However, since $H_{0}$ is a cosmological parameter fitted in the model, the scale factor $a$ is normalized to ensure that the Hubble parameter reaches $H_{0}$ at $a=1$. This is implemented via a rescaling
@@
The power series in the early Universe in Appendix B reveals that the initial conditions for solving the system are functions of the early-Universe value of the torsion scalar field $\varpi$, denoted $\varpi_{r}$. In addition to $\Omega_{L}$$\varpi_{r}$ thus emerges as an independent parameter of the model, playing a crucial role in early-Universe cosmology. The field equations of TorC thus introduce two new parameters, $\Omega_{L}$ and $\varpi_{r}$, which influence the expansion history of the TorC Universe.
@@ II.2 Effective dark energy density and pressure
As discussed in Section I, TorC reproduces the $\Lambda\mathrm{CDM}$ model when $\varpi$ is set to unity. In this limit, the derivatives of $\varpi$ vanish, and the TorC field equations reduce to the Friedmann equations of $\Lambda\mathrm{CDM}$. Therefore, if one wishes to reproduce $\Lambda\mathrm{CDM}$ throughout the whole history of the Universe, one only needs to set the initial value of $\varpi$ to unity, $\varpi_{r}=1$, and the bare dark energy density parameter $\Omega_{L}$ to its correspondent value in $\Lambda\mathrm{CDM}$ model, $\Omega_{L}=\Omega_{\Lambda}$. As $\Lambda\mathrm{CDM}$ emerges as a particular dynamical evolution, based on a choice of initial conditions of the TorC model, TorC can be treated as an extension of $\Lambda\mathrm{CDM}$. The effect of the two additional parameters in TorC can be interpreted as deviations from the dark energy component of $\Lambda\mathrm{CDM}$. Rather than treating dark energy as a constant, it can be redefined as a time-dependent quantity, allowing the effects of the additional TorC terms to be absorbed into an evolving dark energy density and pressure. This reformulation enables the use of Boltzmann codes, such as CAMB, which are capable of incorporating modified dark energy models to investigate the impact of new parameters on the CMB power spectra. In this work, the treatment of dark energy is restricted to the background level, with effects of perturbation assumed to follow the standard $\Lambda\mathrm{CDM}$. The modified first Friedmann equation reflecting a time-dependent dark energy component takes the form
$H^{2}=H_{0}^{2}\bigl{(}\Omega_{r}a^{-4}+\Omega_{m}a^{-3}\bigr{)}+\frac{\rho_{\Lambda}^{\mathrm{eff}}(t)}{3M_{\mathrm{P}}^{2}}.$ (14)
@@
$\rho_{\Lambda}^{\mathrm{eff}}\equiv\frac{3M_{\mathrm{P}}^{2}}{\varpi^{2}}\Biggl{[}$ $H_{0}^{2}\Omega_{L}-H_{0}^{2}\left(\varpi^{2}-1\right)\left(\Omega_{r}a^{-4}+\Omega_{m}a^{-3}\right)$
@@
×$\times\Bigl{(}\Omega_{r}\bigl{(}\varpi^{2}-1\bigr{)}-3a^{4}\Omega_{L}\bigl{(}1+3\varpi^{2}\bigr{)}\Bigr{)}$
@@
$+6\bigl{(}\varpi^{2}-1\bigr{)}\dot{\varpi}\biggl{(}4H_{0}^{2}\bigl{(}\Omega_{r}+\Omega_{m}a+\Omega_{L}a^{4}\bigr{)}$
@@
The effect of $\Omega_{L}$ and $\varpi_{r}$ are thus absorbed into the effective dark energy density and pressure as a modification of time-dependent dark energy. By combining Eqs. 9b and 10, one obtains a coupled system that governs the evolution of the torsion scalar field $\varpi$ and the Hubble parameter. The equation of motion for $\varpi$ determines how the scalar field evolves with time, while the modified Friedmann equation evolves the Hubble parameter, with each depending on the time-dependent values of the other. Together, these form an autonomous and self-consistent dynamical system: given suitable initial conditions, they can be solved simultaneously as a system of ordinary differential equations. This provides the evolution of $\varpi$, its derivative, the Hubble parameter, and the scale factor. These solutions can then be substituted into the expressions for the time-dependent dark energy density and pressure in Eqs. 15 and 17, yielding a complete description of cosmological evolution in the TorC framework. This allows for the exploration of their effects on the expansion history of the Universe and the CMB power spectra.
II.3 Cosmological effects of $\Omega_{L}$ and $\varpi_{r}$
@@
Figure 2: The effect of $\Omega_{L}$ on the comoving Hubble horizon on the left and the CMB power spectrum on the right. The color bar indicates the value of $\Omega_{L}$ in both panels. For both plots, the value for $\varpi_{r}\sim 1$ is set as the $\Lambda\mathrm{CDM}$-equivalent case. The cosmological parameters for computing CMB power spectra are set as the $\Lambda\mathrm{CDM}$ values based on Planck results [30]. For comparison, the case for GR (based on $\Omega_{r}=9.22\times 10^{-5}$$\Omega_{m}=0.314$ and $H_{0}=67.4$ km/s/Mpc) is shown in the plots. A smaller value of $\Omega_{L}$ results in a smaller comoving Hubble horizon at late matter-dominated epochs, increases the amplitude of the CMB power spectrum, and shifts the peaks to lower multipoles.
@@
Figure 3: The effect of $\varpi_{r}$ on the Comoving Hubble Horizon on the left and the CMB power spectrum on the right. The color bar indicates the value of $\varpi_{r}$ in both panels. For both plots, the value for $\Omega_{L}\sim 0.685$ is set as the $\Lambda\mathrm{CDM}$-equivalent case [30]. The other cosmological parameters are set as the values of the $\Lambda\mathrm{CDM}$ model. For comparison, the case for GR (based on $\Omega_{r}=9.22\times 10^{-5}$$\Omega_{m}=0.314$ and $H_{0}=67.4$ km/s/Mpc) is shown in the plots. A smaller value of $\varpi_{r}$ results in a smaller comoving Hubble horizon at early radiation-dominated epoch, increases the amplitude of the CMB power spectrum, and shifts the peaks to higher multipoles.
@@
To investigate the impact of the new TorC parameters on the CMB power spectrum, the model is treated as an extension of $\Lambda\mathrm{CDM}$ and implemented with the CAMB Boltzmann code. CAMB computes the CMB power spectra for a given cosmological model and supports a range of dark energy scenarios by allowing modifications to the dark energy equation of state $w_{\Lambda}^{\mathrm{eff}}\equiv P_{\Lambda}^{\mathrm{eff}}/\rho_{\Lambda}^{\mathrm{eff}}$. However, attempting to incorporate the TorC extension through the equation of state $w_{\Lambda}^{\mathrm{eff}}(a)$ introduces poles when the dark energy density changes sign, as when $\varpi$ exceeds unity in Eq. 15, the second and fourth terms become negative, potentially leading to a negative dark energy density during certain parts of the Universe’s evolution. This obstructs exploration of the full parameter space of $\Omega_{L}$ and $\varpi_{r}$. However, the issue arises only when using $w_{\Lambda}^{\mathrm{eff}}$, and is resolved when pressure and density are implemented separately. To address this, CAMB was therefore modified to evolve the system using the effective dark energy density and pressure directly [^7] Modified CAMB to read external dark energy density and pressure: [37]..
@@
Figure 2 shows the cosmological effects of $\Omega_{L}$. Note that this procesure – adjusting bare dark energy density while holding the other parameters fixed – is not possible in $\Lambda\mathrm{CDM}$ as the density parameters have to sum up to unity. The left panel demonstrates that $\Omega_{L}$ primarily influences the late-time expansion, with its effect most significant at late matter-dominated epochs, where a lower value of $\Omega_{L}$ leads to a smaller comoving Hubble horizon. It’s contribution to the Hubble horizon is negligible at early times, resulting in an evolution of the early Universe that closely resembles $\Lambda\mathrm{CDM}$. The right panel of Fig. 2 shows that a decrease in $\Omega_{L}$ shifts the CMB acoustic peaks to lower multipoles ($\ell$) and enhances their amplitudes. A lower $\Omega_{L}$ causes photon-electron decoupling to occur at higher redshifts. This results in a reduction in both the sound horizon, $r^{*}$, and the angular diameter distance to the last scattering surface, $D_{A}^{*}$. However, a more significant reduction in $D_{A}^{*}$ leads to an increase in the angular scale of the sound horizon, defined as $r^{*}/D_{A}^{*}$, thereby shifting the acoustic peaks toward lower $\ell$ values. Additionally, the earlier decoupling prolongs the period of gravitational potential decay during the radiation-dominated era. This enhanced decay, via the early integrated Sachs-Wolfe (ISW) effect, injects additional energy into the photon-baryon fluid, amplifying the acoustic peak amplitudes.

### v4

@@ I  Introduction
The current cosmological-constant/cold-dark-matter model of cosmology, $\Lambda\mathrm{CDM}$, is built upon two foundations: the theory of general relativity (GR), and the observation that the Universe becomes homogeneous and isotropic at a scale larger than approximately 100 $0\text{\,}\mathrm{Mpc} [1$, 2]. The model has proven success in explaining a wide range of observed phenomena, including the accelerated expansion of the Universe, anisotropy in the cosmic microwave background (CMB) [3], and the observed abundances of deuterium, helium, and other atomic nuclei [4]. Despite its success, the $\Lambda\mathrm{CDM}$ model is currently under scrutiny due to the indications from observational astrophysical data that suggest potential inconsistencies. These include the Hubble tension [5, 6, 7, 8, 9, 10], curvature tension [11, 12], and discrepancies in the light element abundances [13]. Furthermore, the theoretical provenance of dark energy in the form of a very small cosmological constant $\Lambda$, dark matter, and the inflaton field continue to pose a mystery despite decades of research. Another challenge lies in successfully uniting the gravitational force with the other forces of nature described in the standard model (SM) of particle physics. As gravity is non-renormalizable in the framework of quantum field theory, it can unfortunately only be investigated as an effective field theory valid up to the Planck scale [14]. [^1] This is genuinely limiting, because the measured Higgs mass implies runnings such that the SM remains predictive up to this scale. Together, these theoretical challenges and observational tensions motivate the exploration of modified or extended theories of gravity that may offer solutions to these problems [15].
@@
[…]
@@
where $\tensor{\mathcal{R}}{{}_{\mu\nu\sigma\rho}}$ is the curvature, and $\tensor{\mathcal{T}}{{}^{\mu}_{\nu\sigma}}$ is the torsion tensor, with $\tensor{\mathcal{T}}{{}_{\nu}}\equiv\tensor{\mathcal{T}}{{}^{\mu}_{\nu\mu}}$$M_{\mathrm{P}}^{2}=1/8\pi G$ is the Planck mass, $\mu$ and $\nu$ are dimensionless curvature coupling constants; $\lambda$ is a coupling constant of mass dimension two, which can be shown to give rise to a torsional dark energy term intrinsic to the gravitational sector, whereas $\Lambda$ is the usual cosmological constant (which could also be absorbed into the matter Lagrangian $\mathcal{L}_{\mathrm{M}}$). As shown in Appendix A, to ensure the absence of ghost and tachyon degrees of freedom in the tree-level perturbation theory, the following inequality must be satisfied [22]
@@
[…]
@@
In this work, the cosmological implications of the TorC model are investigated, with particular attention given to its potential to alleviate the Hubble tension. TorC introduces two parameters beyond those in $\Lambda\mathrm{CDM}$: the early-Universe value of the torsion scalar field $\varpi_{\mathrm{r}}$ and its velocity, which may be ‘packaged’ into the bare dark energy density parameter $\Omega_{\mathrm{L}}$. To constrain the TorC cosmological parameters, the nested sampling algorithm PolyChord [35] is employed, interfaced via Cobaya [36] with cosmological likelihoods and a modified version of CAMB [37]. A combined analysis is carried out between the Planck 2018 CMB data [34] and the 2020 SH0ES supernova measurement [33], assessing the degree of tension between the two datasets using the $R$-statistic [38, 39, 40].
@@ II  Cosmology
 $\Omega_{\mathrm{L}}$ Bare dark energy density parameter
 $\varpi_{\mathrm{r}}$ Torsion scalar field value in the early Universe
@@
$\Omega_{\mathrm{b}}$ Baryon density parameter
$\Omega_{\mathrm{c}}$ Cold dark matter density parameter
@@
$A_{\mathrm{s}}$ Amplitude of primordial scalar power spectrum
$n_{\mathrm{s}}$ Primordial scalar spectral index
@@
This section focuses on the two additional parameters introduced by the model: $\Omega_{\mathrm{L}}$, the bare dark energy density parameter, and $\varpi_{\mathrm{r}}$, specifying the value of the torsion scalar field in the early Universe, the list of cosmological parameters for TorC is presented in Table 2. In this section, their origin will be outlined: $\Omega_{\mathrm{L}}$ arises from the field equations of TorC, while $\varpi_{\mathrm{r}}$ is set as an initial condition required for solving them. TorC can be treated as an extension of $\Lambda\mathrm{CDM}$ by absorbing the effects of $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$ into an effective dark energy density and pressure. This reformulated dark energy is incorporated into a modified version of CAMB to investigate their impact on the CMB power spectrum.
II.1 TorC cosmological parameters: $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$
@@
$\rho$ $=3M_{\mathrm{P}}^{2}H_{0}^{2}(\Omega_{\mathrm{r}}+\Omega_{\mathrm{m}}+\Omega_{\mathrm{L}}),$ (8a)
$P$ $=3M_{\mathrm{P}}^{2}H_{0}^{2}\left(\frac{\Omega_{\mathrm{r}}}{3}-\Omega_{\mathrm{L}}\right),$ (8b)
where $H_{0}$ is the Hubble parameter today, $\Omega_{\mathrm{r}}$ and $\Omega_{\mathrm{m}}$ are the density parameters of radiation and matter, and $\Omega_{\mathrm{L}}$ represents the ‘intrinsic’ dark energy density parameter, or the ‘bare’ dark energy density parameter.
@@
$H^{2}=$ $H_{0}^{2}\frac{\Omega_{\mathrm{r}}a^{-4}+\Omega_{\mathrm{m}}a^{-3}+\Omega_{\mathrm{L}}}{\varpi^{2}}$
@@
$H^{2}+\dot{H}^{2}=$ $-H_{0}^{2}\frac{\Omega_{\mathrm{r}}a^{-4}+\frac{1}{2}\Omega_{\mathrm{m}}a^{-3}-\Omega_{\mathrm{L}}}{\varpi^{2}}$
@@
$\ddot{\varpi}=$ $\frac{1}{\left(\varpi^{2}-1\right)\left(3\varpi^{2}+1\right)}$
$\times\Bigl{(}-3\varpi\left(\varpi^{2}-1\right)^{2}\left(2H^{2}+\dot{H}\right)$
@@
$H^{2}=H_{0}^{2}\left(\Omega_{\mathrm{r}}a^{-4}+\Omega_{\mathrm{m}}a^{-3}+\Omega_{\Lambda}a^{-4}\right),$ (11)
where $\Omega_{\Lambda}$ is the dark energy density parameter in $\Lambda\mathrm{CDM}$. It is evident that the equations Eq. 9a and Eq. 11 differ in two key ways: first, the density parameters are effectively divided by $\varpi^{2}$ in Eq. 9a; second, an additional term appears in Eq. 9a, involving the scale factor, torsion scalar field, and their derivatives. In standard $\Lambda\mathrm{CDM}$, the dark energy parameter $\Omega_{\Lambda}$ is a constant that can be determined from the density parameters of the other components at $a=1$ and $H=H_{0}$, giving rise to the intuition that the non-gravitational energy budget of the Universe sums to unity when normalized. However, this relationship does not hold for TorC due to the presence of the extra term, similar to how a substantial curvature density would modify the Friedmann equations in GR. As a result, $\Omega_{\mathrm{L}}$ cannot be inferred from the other density parameters: it becomes an independent parameter of the model.
@@
using power series initial conditions for $\varpi$ and $a$ in the early Universe. As the bare dark energy density parameter $\Omega_{\mathrm{L}}$ plays a crucial role in shaping the late-time expansion of the Universe, it leads to a modified value of the current Hubble parameter $H_{0}$. However, since $H_{0}$ is a cosmological parameter fitted in the model, the scale factor $a$ is normalized to ensure that the Hubble parameter reaches $H_{0}$ at $a=1$. This is implemented via a rescaling
@@
The power series in the early Universe in Appendix B reveals that the initial conditions for solving the system are functions of the early-Universe value of the torsion scalar field $\varpi$, denoted $\varpi_{\mathrm{r}}$. In addition to $\Omega_{\mathrm{L}}$$\varpi_{\mathrm{r}}$ thus emerges as an independent parameter of the model, playing a crucial role in early-Universe cosmology. The field equations of TorC thus introduce two new parameters, $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$, which influence the expansion history of the TorC Universe.
@@ II.2 Effective dark energy density and pressure
As discussed in Section I, TorC reproduces the $\Lambda\mathrm{CDM}$ model when $\varpi$ is set to unity. In this limit, the derivatives of $\varpi$ vanish, and the TorC field equations reduce to the Friedmann equations of $\Lambda\mathrm{CDM}$. Therefore, if one wishes to reproduce $\Lambda\mathrm{CDM}$ throughout the whole history of the Universe, one only needs to set the initial value of $\varpi$ to unity, $\varpi_{\mathrm{r}}=1$, and the bare dark energy density parameter $\Omega_{\mathrm{L}}$ to its correspondent value in $\Lambda\mathrm{CDM}$ model, $\Omega_{\mathrm{L}}=\Omega_{\Lambda}$. As $\Lambda\mathrm{CDM}$ emerges as a particular dynamical evolution, based on a choice of initial conditions of the TorC model, TorC can be treated as an extension of $\Lambda\mathrm{CDM}$. The effect of the two additional parameters in TorC can be interpreted as deviations from the dark energy component of $\Lambda\mathrm{CDM}$. Rather than treating dark energy as a constant, it can be redefined as a time-dependent quantity, allowing the effects of the additional TorC terms to be absorbed into an evolving dark energy density and pressure. This reformulation enables the use of Boltzmann codes, such as CAMB, which are capable of incorporating modified dark energy models to investigate the impact of new parameters on the CMB power spectra. In this work, the treatment of dark energy is restricted to the background level, with effects of perturbation assumed to follow the standard $\Lambda\mathrm{CDM}$. The modified first Friedmann equation reflecting a time-dependent dark energy component takes the form
$H^{2}=H_{0}^{2}\bigl{(}\Omega_{\mathrm{r}}a^{-4}+\Omega_{\mathrm{m}}a^{-3}\bigr{)}+\frac{\rho_{\Lambda}^{\mathrm{eff}}(t)}{3M_{\mathrm{P}}^{2}}.$ (14)
@@
$\rho_{\Lambda}^{\mathrm{eff}}\equiv\frac{3M_{\mathrm{P}}^{2}}{\varpi^{2}}\Biggl{[}$ $H_{0}^{2}\Omega_{\mathrm{L}}-H_{0}^{2}\left(\varpi^{2}-1\right)\left(\Omega_{\mathrm{r}}a^{-4}+\Omega_{\mathrm{m}}a^{-3}\right)$
@@
$\times\Bigl{(}\Omega_{\mathrm{r}}\bigl{(}\varpi^{2}-1\bigr{)}-3a^{4}\Omega_{\mathrm{L}}\bigl{(}1+3\varpi^{2}\bigr{)}\Bigr{)}$
@@
$+6\bigl{(}\varpi^{2}-1\bigr{)}\dot{\varpi}\biggl{(}4H_{0}^{2}\bigl{(}\Omega_{\mathrm{r}}+\Omega_{\mathrm{m}}a+\Omega_{\mathrm{L}}a^{4}\bigr{)}$
@@
The effect of $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$ are thus absorbed into the effective dark energy density and pressure as a modification of time-dependent dark energy. By combining Eqs. 9b and 10, one obtains a coupled system that governs the evolution of the torsion scalar field $\varpi$ and the Hubble parameter. The equation of motion for $\varpi$ determines how the scalar field evolves with time, while the modified Friedmann equation evolves the Hubble parameter, with each depending on the time-dependent values of the other. Together, these form an autonomous and self-consistent dynamical system: given suitable initial conditions, they can be solved simultaneously as a system of ordinary differential equations. This provides the evolution of $\varpi$, its derivative, the Hubble parameter, and the scale factor. These solutions can then be substituted into the expressions for the time-dependent dark energy density and pressure in Eqs. 15 and 17, yielding a complete description of cosmological evolution in the TorC framework. This allows for the exploration of their effects on the expansion history of the Universe and the CMB power spectra.
II.3 Cosmological effects of $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$
@@
Figure 2: The effect of $\Omega_{\mathrm{L}}$ on the comoving Hubble horizon on the left and the CMB power spectrum on the right. The color bar indicates the value of $\Omega_{\mathrm{L}}$ in both panels. For both plots, the value for $\varpi_{\mathrm{r}}\sim 1$ is set as the $\Lambda\mathrm{CDM}$-equivalent case. The cosmological parameters for computing CMB power spectra are set as the $\Lambda\mathrm{CDM}$ values based on Planck results [34]. For comparison, the case for GR (based on $\Omega_{\mathrm{r}}=9.22\times 10^{-5}$$\Omega_{\mathrm{m}}=0.314$ and $H_{0}=67.4$ km/s/Mpc) is shown in the plots. A smaller value of $\Omega_{\mathrm{L}}$ results in a smaller comoving Hubble horizon at late matter-dominated epochs, increases the amplitude of the CMB power spectrum, and shifts the peaks to lower multipoles.
@@
Figure 3: The effect of $\varpi_{\mathrm{r}}$ on the Comoving Hubble Horizon on the left and the CMB power spectrum on the right. The color bar indicates the value of $\varpi_{\mathrm{r}}$ in both panels. For both plots, the value for $\Omega_{\mathrm{L}}\sim 0.685$ is set as the $\Lambda\mathrm{CDM}$-equivalent case [34]. The other cosmological parameters are set as the values of the $\Lambda\mathrm{CDM}$ model. For comparison, the case for GR (based on $\Omega_{\mathrm{r}}=9.22\times 10^{-5}$$\Omega_{\mathrm{m}}=0.314$ and $H_{0}=67.4$ km/s/Mpc) is shown in the plots. A smaller value of $\varpi_{\mathrm{r}}$ results in a smaller comoving Hubble horizon at early radiation-dominated epoch, increases the amplitude of the CMB power spectrum, and shifts the peaks to higher multipoles.
@@
To investigate the impact of the new TorC parameters on the CMB power spectrum, the model is treated as an extension of $\Lambda\mathrm{CDM}$ and implemented with the CAMB Boltzmann code. CAMB computes the CMB power spectra for a given cosmological model and supports a range of dark energy scenarios by allowing modifications to the dark energy equation of state $w_{\Lambda}^{\mathrm{eff}}\equiv P_{\Lambda}^{\mathrm{eff}}/\rho_{\Lambda}^{\mathrm{eff}}$. However, attempting to incorporate the TorC extension through the equation of state $w_{\Lambda}^{\mathrm{eff}}(a)$ introduces poles when the dark energy density changes sign, as when $\varpi$ exceeds unity in Eq. 15, the second and fourth terms become negative, potentially leading to a negative dark energy density during certain parts of the Universe’s evolution. This obstructs exploration of the full parameter space of $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$. However, the issue arises only when using $w_{\Lambda}^{\mathrm{eff}}$, and is resolved when pressure and density are implemented separately. To address this, CAMB was modified to evolve the system using the effective dark energy density and pressure directly. [^7] Modified CAMB to read external dark energy density and pressure: [41].
@@
Fig. 2 shows the cosmological effects of $\Omega_{\mathrm{L}}$. Note that this procesure – adjusting bare dark energy density while holding the other parameters fixed – is not possible in $\Lambda\mathrm{CDM}$ as the density parameters have to sum up to unity. The left panel demonstrates that $\Omega_{\mathrm{L}}$ primarily influences the late-time expansion, with its effect most significant at late matter-dominated epochs, where a lower value of $\Omega_{\mathrm{L}}$ leads to a smaller comoving Hubble horizon. It’s contribution to the Hubble horizon is negligible at early times, resulting in an evolution of the early Universe that closely resembles $\Lambda\mathrm{CDM}$. The right panel of Fig. 2 shows that a decrease in $\Omega_{\mathrm{L}}$ shifts the CMB acoustic peaks to lower multipoles ($\ell$) and enhances their amplitudes. A lower $\Omega_{\mathrm{L}}$ causes photon-electron decoupling to occur at higher redshifts. This results in a reduction in both the sound horizon, $r^{*}$, and the angular diameter distance to the last scattering surface, $D_{A}^{*}$. However, a more significant reduction in $D_{A}^{*}$ leads to an increase in the angular scale of the sound horizon, defined as $r^{*}/D_{A}^{*}$, thereby shifting the acoustic peaks toward lower $\ell$ values. Additionally, the earlier decoupling prolongs the period of gravitational potential decay during the radiation-dominated era. This enhanced decay, via the early integrated Sachs-Wolfe (ISW) effect, injects additional energy into the photon-baryon fluid, amplifying the acoustic peak amplitudes.